OpenStudy (anonymous):

Use logarithmic differentiation to find the derivative of the function. y=x1/x

OpenStudy (anonymous):

I got \[(x^1/x)(1/x^2lnx+1/x^2)\] this is not right what did i do wrong

OpenStudy (anonymous):

[multiply every variable by xy] y + x = xy y - xy = - x y(1 - x) = - x y = - [x / (1 - x)] y = - x / (1 - x)

zepdrix (zepdrix):

 Hmm your answer looks good, I think you just missed a negative sign in one spot.